# SVP: Style-Enhanced Vivid Portrait Talking Head Diffusion Model

### v1

@@ Introduction
V-Express 340.01 33.66 27.51 0.8939 5.28 8.12 125.36 12.33 26.35 0.8688 30.58 35.56
Hallo 445.49 35.06 25.84 0.8666 9.60 12.42 231.96 15.31 22.44 0.8078 31.86 35.63
AniPortrait 460.04 40.65 27.16 0.8982 5.86 9.03 292.18 13.27 25.27 0.8616 31.20 36.37
SadTalker 492.4 52.94 27.33 0.8863 7.23 11.03 536.68 23.7 22.62 0.8042 28.81 32.60
Ours 235.70 28.80 28.87 0.9126 4.64 6.45 102.40 10.39 26.38 0.8800 22.95 26.01
@@
V-Express 367.91 63.17 21.23 0.7904 7.41 9.33
Hallo 399.18 59.05 19.70 0.7573 26.16 28.54
AniPortrait 535.53 68.18 19.34 0.7583 26.55 29.81
SadTalker 492.83 82.71 20.07 0.7670 27.00 29.55
Ours w/o style 485.47 57.96 21.73 0.8151 8.49 10.20
Ours 392.31 56.69 22.07 0.8210 6.76 8.44
@@
Experiments Setting
$\mathtt{SVP}$ is implemented using PyTorch (Paszke et al. 2019) and optimized with Adam (Kingma and Ba 2014). The intrinsic style encoder is trained on the MEAD (Wang et al. 2020) and HDTF (Zhang et al. 2021) datasets. During training, we consider samples with the same identity and emotion in MEAD as positive samples, and segments from the same video in HDTF as positive samples. Additionally, we will randomly dropout expression coefficients or audios, but they will not be zeroed out simultaneously.
@@
We compare our method with several previous works, including SadTalker (Zhang et al. 2023), AniPortrait (Wei, Yang, and Wang 2024), V-Express (Wang et al. 2024), and Hallo (Xu et al. 2024).
@@
To demonstrate the superiority of the proposed method, we evaluate the model using several quantitative metrics. We utilize the Fréchet Inception Distance (FID) (Heusel et al. 2017) to assess the quality of the generated frames, and further employe the Fréchet Video Distance (FVD) (Unterthiner et al. 2018) for video-level evaluation. To evaluate the quality of the generated talking head videos, Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) are adopted. To evaluate lip-sync accuracy, we use the Mouth Landmark Distance (M-LMD) (Bulat and Tzimiropoulos 2017). For assessing the accuracy of the generated facial expressions, we use the Full-Face Landmark Distance (F-LMD).
@@
As shown in Table 1, in the video reconstruction experiments, our method achieves the best performance across all the metrics on both MEAD and HDTF datasets. We have a significant advantage in evaluating the quality of video and single-frame images, as evidenced by the lower FVD and FID scores. The SSIM and PSNR scores indicate that the quality of the videos reconstructed by our method is significantly better than that of other methods. The M-LMD scores demonstrate that our method achieves more accurate lip-sync, while the F-LMD scores reflect that our method better restores facial expressions through intrinsic style. These results indicate that using intrinsic style can significantly enhance the quality of video generation.
@@
To demonstrate the intrinsic style transfer capability of our method, we conducted expression transfer experiments in addition to video reconstruction experiments. This experiment can only be performed on the MEAD dataset, which contains multiple emotions for a single identity. Specifically, we used the neutral expression faces from the dataset as references and used videos with distinct expressions (such as happy, sad, and angry) to drive the generation. In such an experimental setup, methods that rely solely on audio for driving expressions struggle to effectively convey the corresponding emotions. In this scenario, our method still maintains its superiority. As shown in Table 2, although the results of all methods deteriorate compared to the reconstruction experiments, $\mathtt{SVP}$ still leads in all metrics.
Qualitative Comparison
@@
In Figure 3, we present a comparison of the visual effects between ours and other methods. The first two rows show the comparison of reconstruction results, while the last two rows show the comparison of intrinsic style transfer results. In the reconstruction experiments, our method achieves accurate synchronization of head movements, lip shapes, and even eye blinking, while effectively preserving the identity of the speaker. In the intrinsic style transfer experiments, when there are significant differences in expressions between the reference face and the real video, our method effectively transfers the expressions and details of the face in the style reference video, while maintaining consistency in other conditions.
Ablation Study
@@
In diffusion models, if we do not employ a probabilistic learning and instead use a deterministic style prior for training, it may lead to overfitting of the training results. Figure 4 shows the results of different intrinsic style acquisition methods. Using a deterministic intrinsic style causes the model to transfer (eye reflections/facial contours) from the style reference video to the new face, leading to issues with identity deviation. By employing the probabilistic learning and resampling method, the intrinsic style prior obtained by the model from the same training video varies each time, thereby preventing the transfer of incorrect content to the generated video.
@@
We use t-distributed Stochastic Neighbor Embedding (t-SNE) (Van der Maaten and Hinton 2008) to project the intrinsic style priors into a two-dimensional space. Figure 5 shows the intrinsic style priors of a speaker from the MEAD dataset. Each code is color-coded according to its corresponding emotion and intensity. The style priors with the same emotion first cluster together. Within each cluster, the style priors with the same intensity are closer to each other, and there are noticeable transitions between intrinsic style priors of different intensities. These observations indicate that our model can learn a continuous distribution of intrinsic styles. As shown in Figure  6, when performing linear interpolation between two intrinsic style priors extracted from the test set, the facial expressions and details in the generated video transition smoothly.
Conclusion

### v2

@@ 4 Experiments
V-Express 340.01 33.66 27.51 0.8939 5.28 8.12 2.10/4.59 125.36 12.33 26.35 0.8688 30.58 35.56 2.70/6.05
Hallo 445.49 35.06 25.84 0.8666 9.60 12.42 1.70/5.94 231.96 15.31 22.44 0.8078 31.86 35.63 2.10/6.74
AniPortrait 460.04 40.65 27.16 0.8982 5.86 9.03 8.30/1.40 292.18 13.27 25.27 0.8616 31.20 36.37 4.00/0.53
EAMM 529.34 58.77 21.88 0.8139 11.88 12.43 1.90/3.98 797.23 43.46 18.47 0.7202 29.93 30.48 2.40/4.54
SadTalker 492.40 52.94 27.33 0.8863 7.23 11.03 0.40/4.74 536.68 23.7 22.62 0.8042 28.81 32.60 0.50/5.96
Ours 235.70 28.80 28.87 0.9126 4.64 6.45 0.30/4.76 102.40 10.39 26.38 0.8800 22.95 26.01 0.20/5.40
@@
V-Express 367.91 63.17 21.23 0.7904 7.41 9.33 2.03/4.25 0.7819
Hallo 399.18 59.05 19.70 0.7573 26.16 28.54 1.93/4.85 0.7851
AniPortrait 535.53 68.18 19.34 0.7583 26.55 29.81 8.67/1.32 0.7711
EAMM 550.90 63.94 21.25 0.7848 12.51 12.72 2.17/4.20 0.7349
SadTalker 492.83 82.71 20.07 0.7670 27.00 29.55 0.70/4.93 0.7417
Ours w/o style 485.47 57.96 21.73 0.8151 8.49 10.20 0.93/4.03 0.7879
Ours 392.31 56.69 22.07 0.8210 6.76 8.44 0.77/4.14 0.8473
@@
4.1 Experiments Setting
$\mathtt{SVP}$ is implemented using PyTorch [14] and optimized with Adam [11]. The intrinsic style encoder is trained on the MEAD [27] and HDTF [37] datasets. During training, we consider samples with the same identity and emotion in MEAD as positive samples, and segments from the same video in HDTF as positive samples. Additionally, we will randomly dropout expression coefficients or audios, but they will not be zeroed out simultaneously.
@@ 4.2 Quantitative Comparison
We compare our method with several previous works, including EAMM [10], SadTalker [36], AniPortrait [30], V-Express [25], and Hallo [34].
@@
To demonstrate the superiority of the proposed method, we evaluate the model using several quantitative metrics. We utilize the Fréchet Inception Distance (FID) [8] to assess the quality of the generated frames, and further employe the Fréchet Video Distance (FVD) [22] for video-level evaluation. To evaluate the quality of the generated talking head videos, Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) are adopted. To evaluate lip-sync accuracy, we use the Mouth Landmark Distance (M-LMD) [3] and the average visual-audio offset and confidence of SyncNet [6]. For assessing the accuracy of the generated facial expressions, we use the Full-Face Landmark Distance (F-LMD).
Additionally, we introduce Intrinsic Style Similarity (StyleSim) to evaluate the performance of the generated results in terms of facial expressions and details. For the generated video $V_{res}$ and $V_{gt}$, we use the 3DMM encoder to extract their sequential expression parameters $\beta_{res}$ and $\beta_{gt}$. Then, we use the pretrained style encoder from StyleTalk [13] to encode them into style representations $s_{res}$ and $s_{gt}$. We consider the cosine similarity between $s_{res}$ and $s_{gt}$ as StyleSim.
@@
As shown in Table 1, in the video reconstruction experiments, our method achieves the best performance in most metrics on both MEAD and HDTF datasets. We have a significant advantage in evaluating the quality of video and single-frame images, as evidenced by the lower FVD and FID scores. The SSIM and PSNR scores indicate that the quality of the videos reconstructed by our method is significantly better than that of other methods. The M-LMD and SyncNet-offset scores demonstrate that our method achieves more accurate lip-sync, while the F-LMD scores reflect that our method better restores facial expressions through intrinsic style. These results indicate that using intrinsic style can significantly enhance the quality of video generation.
@@
To demonstrate the intrinsic style transfer capability of our method, we conducted expression transfer experiments in addition to video reconstruction experiments. This experiment can only be performed on the MEAD dataset, which contains multiple emotions for a single identity. Specifically, we used the neutral expression faces from the dataset as references and used videos with distinct expressions (such as happy, sad, and angry) to drive the generation. In such an experimental setup, methods that rely solely on audio for driving expressions struggle to effectively convey the corresponding emotions. As shown in Table 2, $\mathtt{SVP}$ still leads in most metrics and maintains its superiority in this scenario. For the additionally calculated StyleSim, our method significantly outperforms other methods with intrinsic style, indicating that the transfer of intrinsic style can better control facial expressions and details.
4.3 Qualitative Comparison
@@
In Figure 3(a), we present a visual comparison of our method with other methods, including comparisons of reconstruction results and intrinsic style transfer results. In the reconstruction experiments, our method achieves accurate synchronization of head movements, lip shapes, and even eye blinking, while effectively preserving the identity of the speaker. In the intrinsic style transfer experiments, when there are significant differences in expressions between the reference face and the real video, our method effectively transfers the expressions and details of the face in the style reference video, while maintaining consistency in other conditions. In Figure 3(b), we show the generated results of our method on different types of portraits. Our method successfully generates videos with rich expressions and natural movements on out-of-domain data, even non-human portraits, demonstrating strong robustness.
4.4 Ablation Study
@@
When training the style layer of the diffusion model in Sec. 3.2, if we do not employ a probabilistic learning and instead use a deterministic style prior for training, it may lead to overfitting of the training results and lost identity information. Figure 4 shows the results of different intrinsic style acquisition methods. Using a deterministic intrinsic style causes the model to transfer (eye reflections/facial contours) from the style reference video to the new face, leading to issues with identity deviation. By employing the probabilistic learning and resampling method, the intrinsic style prior obtained by the model from the same training video varies each time, thereby preventing the transfer of incorrect content to the generated video.
@@
We use t-distributed Stochastic Neighbor Embedding (t-SNE) [23] to project the intrinsic style priors into a two-dimensional space. Figure 5 shows the intrinsic style priors of a speaker from the MEAD dataset. Each code is color-coded according to its corresponding emotion and intensity. The style priors with the same emotion first cluster together. Within each cluster, the style priors with the same intensity are closer to each other, and there are noticeable transitions between intrinsic style priors of different intensities. These observations indicate that our model can learn a continuous distribution of intrinsic styles. As shown in Figure  6, when performing linear interpolation between two intrinsic style priors extracted from the test set, the facial expressions and details in the generated video transition smoothly.
5 Conclusion